# Autoregressive Image Generation with Randomized Parallel Decoding

### v2

@@ 3.1 Rethinking Visual Autoregressive Modeling
Figure 3: Analysis of attention scores. Normalized attention maps from select heads in the final layer of RandAR (Pang et al., 2025). The maps, partitioned by token type (masked vs. unmasked), reveal that attention weights are predominantly concentrated on unmasked tokens.
@@ 4.1 Implementation Details
We train the class-conditional model on ImageNet-1K ($\times256$) (Deng et al., 2009) for 400 epochs and the text-to-image model on a 4M subset of BLIP-3o (Chen et al., 2025) for 50 epochs. All models are optimized with the AdamW (Loshchilov & Hutter, 2019) optimizer using a learning rate of 1e-4 per 256 batch size. Further configuration details are in Appendix A.
@@
We use FID (Heusel et al., 2017) as the primary evaluation metric, and additionally report Inception Score (IS) (Salimans et al., 2016), precision, and recall (Kynkäänniemi et al., 2019), following the protocol of ADM (Dhariwal & Nichol, 2021). For text-to-image generation, we evaluate our models using selected metrics from GenEval (Ghosh et al., 2023). The efficiency profile is evaluated on an NVIDIA A800-80GB GPU, considering only the token generation process.
@@
Diffusion DiT-L/2 (Peebles & Xie, 2023) 458 M 250 1.32 1.62 5.02 167.2 0.75 0.57
DiT-XL/2 (Peebles & Xie, 2023) 675 M 250 0.91 2.14 2.27 278.2 0.83 0.57
@@ 4.2 Image Generation
Table 3: Quantitative evaluation of text-to-image generation.
@@ Appendix A Implementation Details
We detail the hyperparameters for the experimental setups to ensure the precise reproducibility of our findings. The AdamW (Kingma, 2014; Loshchilov & Hutter, 2019) optimizer, with a consistent configuration, was utilized in all experiments. The batch size and learning rate were tailored for each model based on computational resources.
@@
For the sampling process, all models employ an arccos schedule to determine the number of tokens to decode at each step. A linear scheduler is utilized for the classifier-free guidance (CFG) (Ho & Salimans, 2022) scale. Optimal performance is achieved by setting both the temperature and top-p to 1.0, while top-k sampling is not applied (denoted as None in the Tab. 6, which is equivalent to setting top-k to the vocabulary size of 16,384).
@@ Appendix B Additional Experiments
We also evaluated the performance of ARPG under specific orders, including raster order and several alternatives (Esser et al., 2021), as shown in Tab. 7. While random-order modeling is more challenging due to the $n!$ possible orderings, it still outperforms generation with fixed orders. The constraint of a fixed order impedes effective parallel decoding and zero-shot tasks.
@@ Appendix E Pseudo-Code
# shuffle input
@@
targets = shuffled_ids.clone().detach()
# Shuffle RoPE using the same order.
@@
# compute loss
return cross_entropy(logits, targets)

### v3

@@ 1 Introduction
Furthermore, we extend our method to more complex tasks, including controllable generation, zero-shot generalization, and text-to-image generation, as shown in Fig. 1. In summary:
@@ 3.1 Rethinking Visual Autoregressive Modeling
Figure 3: Analysis of attention scores. Normalized attention maps from multiple distinct heads in the final layer of RandAR (Pang et al., 2025). The maps, partitioned by token type (masked vs. unmasked), reveal that attention weights are predominantly concentrated on unmasked tokens.
@@ 4.1 Implementation Details
We train the class-conditional model on ImageNet-1K ($\times256$) (Deng et al., 2009) for 400 epochs and the text-to-image model on a 4M subset of BLIP-3o (Chen et al., 2025) for 50 epochs. All models are optimized with the AdamW (Loshchilov and Hutter, 2019) optimizer using a learning rate of 1e-4 per 256 batch size. Further configuration details are in Appendix A.
@@
We use FID (Heusel et al., 2017) as the primary evaluation metric, and additionally report Inception Score (IS) (Salimans et al., 2016), precision, and recall (Kynkäänniemi et al., 2019), following the protocol of ADM (Dhariwal and Nichol, 2021). For text-to-image generation, we evaluate our models using selected metrics from GenEval (Ghosh et al., 2023). The efficiency profile is evaluated on an NVIDIA A800-80GB GPU, considering only the token generation process.
@@
Diffusion DiT-L/2 (Peebles and Xie, 2023) 458 M 250 1.32 1.62 5.02 167.2 0.75 0.57
DiT-XL/2 (Peebles and Xie, 2023) 675 M 250 0.91 2.14 2.27 278.2 0.83 0.57
@@ 4.2 Image Generation
Table 3: Quantitative evaluation of text-to-image generation at $\times512$ resolution.
@@ 5 Conclusion
Acknowledgments
This work is supported by the Young Scientists Fund of the National Natural Science Foundation of China (NSFC) (No. 62506305), the Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), the Key Research and Development Program of Zhejiang Province (No. 2025C01026), the Scientific Research Project of Westlake University (No. WU2025WF003), and the Chinese Association for Artificial Intelligence (CAAI) & Ant Group Research Fund - AGI Track (No. 2025CAAI-ANT-13). It is also supported by the research funds of the National Talent Program and the Hangzhou Municipal Talent Program.
Additionally, this work is partially supported by the CAS Project for Young Scientists in Basic Research (YSBR-116), the National Natural Science Foundation of China (NSFC) (No. 62325603, No. 62236009, No. U22A20103), the Beijing Science and Technology Plan (No. Z241100004224011), and Shanghai NeuHelium Neuromorphic Technology Co., Ltd.
@@ Appendix A Implementation Details
We detail the hyperparameters for the experimental setups to ensure the precise reproducibility of our findings. The AdamW (Kingma, 2014; Loshchilov and Hutter, 2019) optimizer, with a consistent configuration, was utilized in all experiments. The batch size and learning rate were tailored for each model based on computational resources.
@@
For the sampling process, all models employ an arccos schedule to determine the number of tokens to decode at each step. A linear scheduler is utilized for the classifier-free guidance (CFG) (Ho and Salimans, 2022) scale. Optimal performance is achieved by setting both the temperature and top-p to 1.0, while top-k sampling is not applied (denoted as None in the Tab. 6, which is equivalent to setting top-k to the vocabulary size of 16,384).
@@ Appendix B Additional Experiments
We also evaluated the performance of ARPG under specific orders, including raster order and several alternatives (Esser et al., 2021), as shown in Tab. 7. While random-order modeling is more challenging due to the $n!$ possible orderings, it still outperforms fixed orders. The constraint of a fixed order impedes effective parallel decoding and zero-shot tasks.
@@ Appendix E Pseudo-Code
# shuffle input and RoPE using the same order.
@@
return cross_entropy(logits, shuffled_ids.clone().detach())
@@
Appendix F Discussion
In future work, we plan to extend ARPG to large-scale text-to-image synthesis and unified models for understanding and generation (Xie et al., 2024; 2025). Additionally, we will explore architectures and methods that combine the strengths of AR and diffusion models (Yu et al., 2025; Chen et al., 2025; Gao and Shou, 2025), incorporating techniques such as feature caching Ma et al. (2024); Liu et al. (2025), quantization Tao et al. (2025); Li* et al. (2025) and pruning Zhu et al. (2025); Feng et al. (2024) to further enhance efficiency.
Appendix G The Use of Large Language Models (LLMs)
Our use of LLMs was strictly limited to proofreading for grammatical and spelling errors. LLMs made no contribution to the ideation process or the writing of the content.